OpenStudy (anonymous):

**MEDAL** Use a clock or stopwatch to take your pulse rate while sitting relaxed. Record the number of beats for thirty seconds. Now, walk out your front door and run around your house, then come back inside. Take your pulse again and record the number of beats for thirty seconds. What is the percent increase in your pulse rate? Show your work. Round your percent to the nearest whole number.

OpenStudy (anonymous):

Ok you ready?

OpenStudy (anonymous):

Say your rest heart beat was 35 beats over 30 seconds - This gives 70 Beats / Minute

OpenStudy (anonymous):

And say after you run it was 60 beats in 30 seconds - This gives 120 Beats / Minute

OpenStudy (anonymous):

So we have Measure 1 (M1 = 70b/m) Measure 2 (M2 = 120b/m) Are you with me so far at how we got these values? M1=70bpm M2=120bpm

OpenStudy (anonymous):

So if you look at the difference between the two pulses 120-70(bpm), we can see how much our pulse increased in bpm. 120-70 = 50

OpenStudy (anonymous):

70 beats/60 seconds is 7/6 (1.17) 120 beats/60 seconds is 12/6 (2) 2-1.17 = 0.83... and the difference of pulse rate was 50 0.83 * 50 = 41.5 or 41.5% Answer should be 41.5% and sorry I didn't explain it better lol

OpenStudy (anonymous):

thankss!